# Quasi-alternating surgeries

### v1

@@ 7. The quasi-alternating slopes of torus knots
Here we classify the quasi-alternating slopes of torus knots. For coprime integers $a>b>1$, we denote by $T_{a,b}$ the $(a,b)$-torus knot.
@@
Since a Dehn surgery on a torus knot is either a lens space, a connected sum of lens spaces, or a small Seifert fibered space [Mos71] and lens spaces are double branched covers of $2$-bridge links (which are alternating), we only need to understand which small Seifert fibered spaces are double branched covers of quasi-alternating links. For that, we appeal to Issa’s work on quasi-alternating Montesinos links [Iss18] and Moser’s classification of surgeries on torus knots [Mos71].
Theorem 7.2 (Moser [Mos71])
@@
with surgery diagram of Figure 2 where $c$ and $d$ are integers such that $ac+bd=-1$. If $p/q\neq ab$, then this is a Seifert fibered space. If $p/q=ab$ then it is a generalized Seifert fibered space diffeomorphic to $L(a,b)\#L(b,a)$.
Recall that we use Regina’s notation for Seifert fibered spaces [BBPea]. Since the notation and conventions used in [Mos71] differ from ours, we briefly and roughly sketch a proof using rational Kirby calculus.
@@
00
Figure 2. A surgery diagram of the Seifert fibered space $SFS[S2:(a,d)(b,c)(abq-p,q)]$
@@
Figure 2 shows a surgery diagram for these manifolds. By a slam dunk on the knot with surgery coefficient $a/d$ into the 00 framed component, we obtain a knot $K_{0}$ with surgery coefficient $-d/a$ linked by the components $K_{1}$ and $K$ with surgery coefficients $b/c$ and $ab-p/q$ respectively. Together $K_{0}\cup K_{1}$ form a Hopf link where $K$ is a curve in the Heegaard torus between them that is isotopic to the meridian of $K_{0}$ and the longitude of $K_{1}$. Since $ac+bd=-1$, a sequence of Rolfsen twists on $K_{0}$ and $K_{1}$ informed from a continued fraction expansion of $-d/a$ with partial fraction $b/c$ reduces both the surgery coefficients of $K_{0}$ and $K_{1}$ to $1/0$, twists $K$ to the $(a,b)$ curve on the Heegaard torus, and reduces its surgery coefficient to $p/q$. ∎
In [Iss18] Issa classifies quasi-alternating Montesinos knots as follows.
Theorem 7.3
@@
    $e=n-1$ and $\frac{\beta_{i}}{\alpha_{i}}+\frac{\beta_{j}}{\alpha_{j}}<1$ for some $i\neq j$.
In addition, it is shown in [Iss18, Corollary 1] that a Seifert fibered space is the double branched cover of a quasi-alternating link if and only if it is the double branched cover of a quasi-alternating Montesinos link. The Seifert fibered space $SFS[S2:(\alpha_{1},\beta_{1})\ldots(\alpha_{n},\beta_{n})]$ in Regina’s notation is the double branched cover of the Montesinos knot $M(0;\alpha_{i}/\beta_{i},\ldots,\alpha_{n}/\beta_{n})$ which is in general not in Issa’s standard form.
@@
Now by Theorem 7.2, $(p/q)$-surgery on $T_{a,b}$ is diffeomorphic to the (generalized) Seifert fiber space
@@
where $c$ and $d$ are any integers such that $ac+bd=-1$. Since $a>b>0$ it follows that $d>0>c$. Furthermore the integers $c,d$ satisfying $ac+bd=-1$ may be chosen so that $a>d>0$ and $b>-c$, and hence $\frac{a}{d}>1$ and $\frac{b}{b+c}>1$.
Also, notice that if $p/q=ab-1$ then the above manifold is a lens space, and thus the slope $ab-1$ is quasi-alternating. From [Wat11] it follows that any slope $p/q\geq ab-1$ is quasi-alternating. Otherwise, we have $p/q<ab-1$ so that $\frac{abq-p}{q}>1$ and Theorem 7.3 informs us that the Montesinos link $M\left(1;a/d,b/(b+c),ab-p/q\right)$ is quasi-alternating if and only if one of the following is fulfilled:
@@
(1) cannot occur because $ad+bc=-1$. Either (2) or (3) occurs if and only if either $\frac{q}{abq-p}>\frac{-c}{b}$ or $\frac{q}{abq-p}>\frac{a-d}{a}$. Since $abq-p>q>0$, $a>d>0$, and $b>-c>0$, this is occurs if and only if $\frac{p}{q}>ab-\max\{\frac{a}{a-d},\frac{b}{-c}\}$. Thus, with [Iss18, Corollary 1], this gives the claimed classification of quasi-alternating slopes once we set $m=a-d$ and $n=-c$. ∎
@@ 8. Formal L-space surgeries
In this section, we prove the results about formal L-space surgeries.
@@
Let $K$ be a knot in $S^{3}$ and let $r=p/q\in\mathbb{Q}$ be slope with $q\geq 1$. As illustrated in Figure 3, the manifold $K(p/q)$ is contained in a triad consisting of $K(p/q)$, $U(q/r)$ and $K((p+q)/q)$, where $p/q=n-r/q$ with $n\in\mathbb{Z}$ and $q>r\geq 0$. Since every lens space is a formal L-space, we see that $K((p+q)/q)$ is a formal L-space, whenever $K(p/q)$ is a formal $L$-space and $p/q>0$. ∎
@@
Let $K$ be any asymmetric knot admitting an alternating or quasi-alternating surgery. For example, let $K$ be one of the asymmetric $L$-space knots in the SnapPy census or one of the Baker–Luecke asymmetric $L$-space knots. For all $r$ sufficiently large $K(r)$ is an asymmetric L-space. However, infinitely many of these surgeries are also formal L-spaces. ∎
@@
where $a_{i}\geq 2$ for all $i$. We call $\ell$ the length of this continued fraction expansion. Note also that if $p/q>n$, then $a_{1}>n$. We prove the theorem by induction on $\ell$ and then by induction on the coefficient $a_{\ell}$. The base case for this induction is the case $\ell=1$. In this case, $p/q$ is an integer and Theorem 8.1 gives the necessary result. Thus suppose that
@@
has continued fraction expansion of length $\ell>1$.
Consider also the rational numbers
$p_{0}/q_{0}=[a_{1},\dots,a_{\ell-1}]^{-}$
and
$p_{1}/q_{1}=[a_{1},\dots,a_{\ell}-1]^{-}.$
Figure 4 shows that the manifolds $K(p_{0}/q_{0})$, $K(p_{1}/q_{1})$ and $K(p/q)$ sit in a surgery triad. Furthermore standard continued fraction identities show that $\frac{p}{q}=\frac{p_{0}+p_{1}}{q_{1}+q_{0}}$. Thus we see that if $K(p_{0}/q_{0})$ and $K(p_{1}/q_{1})$ are formal L-spaces, then $K(p/q)$ is also a formal L-space. Using this we may induct on $a_{\ell}\geq 2$. Note that $K(p_{0}/q_{0})$ is a formal L-space by the induction on $\ell$. If $a_{\ell}=2$, then let $\ell^{\prime}\geq 1$ be maximal such that $a_{\ell^{\prime}}>2$. We then have that
$p_{1}/q_{1}$ $=[a_{1},\dots,a_{\ell^{\prime}},2,\dots,2,1]^{-}$
$=[a_{1},\dots,a_{\ell^{\prime}}-1]^{-}$
That is $p_{1}/q_{1}$ has continued fraction of length $\ell^{\prime}<\ell$ and so we may conclude that $K(p_{1}/q_{1})$ is a formal L-space. If $a_{\ell}>2$, then the inductive hypothesis implies that $K(p_{1}/q_{1})$ is a formal L-space. In either case we may deduce that $K(p/q)$ is also a formal L-space. ∎
@@
Surgery on a torus knot is a Seifert fibered space over $S^{2}$ (including the possibility of a lens space) or a connected sum of two lens spaces. In his classification of quasi-alternating Montesinos knots, Issa showed that a Seifert fibered space over $S^{2}$ is a formal L-space if and only if it is the double branched cover of a quasi-alternating link [Iss18]. ∎
@@
A formal L-space is the boundary of a negative-definite manifold with $H_{1}=0$. If $K(r)$ bounds a negative definite manifold with $H_{1}=0$, then $K(\lceil r\rceil)$ also bounds such a manifold. The work of Greene implies that we then have the following bound [Gre15].
$2g(K)\leq\lceil r\rceil-\sqrt{\lceil r\rceil}.$
This can be rearranged to obtain a lower bound for $\lceil r\rceil$ which is then converted to a lower bound for $r$ using the fact that $r>\lceil r\rceil-1$. ∎
@@
Figure 3. The triad obtained by performing 00, $1/0$ and $-1$ surgery on the knot $C$
[…]
Figure 4. The triad obtained by performing 00, $1/0$ and $-1$ surgery respectively on the knot $C$ yields $K(p_{0}/q_{0})$, $K(p_{1}/q_{1})$ and $K(p/q)$.
@@ 9. Observations and questions
While the strategies from Section 3 together with the main result of [CK09] and [Wat17] work well to determine the quasi-alternating/thin status of large infinite subsets of slopes for a given knot, in general this does not give a full classification of quasi-alternating/thin slopes.
@@
Consider the torus knot $T_{5,3}$. Theorem 7.1 shows that the quasi-alternating slopes are $(12+1/2,\infty)$. In Example 1.1 we have observed that the $(12+1/2)$-surgery on $T_{5,3}$ is the double branched cover of the knot $K11n50$ which is thin but not quasi-alternating [Gre10]. If we apply [Wat17] we get that all slopes smaller or equal to $12$ are thick. Thus the slopes in the interval $(12,12+1/2)$ remain of undetermined thick/thin status. However, experiments suggest that all surgeries in that interval are thin. For example: $K10n21(12+1/3)=DBC(K12n296)$ and $K10n21(12+1/4)=DBC(K13n2006)$ and both branching sets are thin but not quasi-alternating.
@@
What is the classification of the thin slope of torus knots?
Most surgeries on torus knots are small Seifert fibered spaces [Mos71] and these arise as double branched covers of Montesinos links and Seifert links. Since the Seifert links with L-space double branched covers are also Montesinos links [BGH24], this naturally leads to the question of thinness for Montesinos links.
@@
Which Montesinos links are thin?

### v2

@@ 7. The quasi-alternating slopes of torus knots
In the previous sections, we dealt with finite families of hyperbolic knots. In this section, we address an infinite family of non-hyperbolic knots, the collection of torus knots. We classify the quasi-alternating slopes of torus knots. For coprime integers $a>b>1$, we denote by $T_{a,b}$ the $(a,b)$-torus knot.
@@
A slope $p/q\in\mathbb{Q}$ is quasi-alternating for the torus knot $T_{a,b}$ with $a>b>1$ if and only if
@@
Since a Dehn surgery on a torus knot is either a lens space, a connected sum of lens spaces, or a small Seifert fibered space [Moser1971elementary] and lens spaces are double branched covers of $2$-bridge links (which are alternating), we only need to understand which small Seifert fibered spaces are double branched covers of quasi-alternating links. For that, we appeal to Issa’s work on quasi-alternating Montesinos links [Is18] and Moser’s classification of surgeries on torus knots [Moser1971elementary].
Theorem 7.2 (Moser [Moser1971elementary])
@@
with surgery diagram of Figure 2 where $c$ and $d$ are integers such that ac+bd=1ac+bd=-1. If $p/q\neq ab$, then this is a Seifert fibered space. If $p/q=ab$ then it is a generalized Seifert fibered space diffeomorphic to L(a,b)#L(b,a)L(a,b)\#L(b,a).
Recall that we use Regina’s notation for Seifert fibered spaces [BBP+]. Since the notation and conventions used in [Moser1971elementary] differ from ours, we briefly and roughly sketch a proof using rational Kirby calculus. Essentially the same proof can also be found in Lemma 4.4 of [Owens_Strle12].
@@
0
Figure 2. A surgery diagram of the Seifert fibered space SFS[S2:(a,d)(b,c)(abqp,q)]SFS[S2:(a,d)(b,c)(abq-p,q)]
@@
Figure 2 shows a surgery diagram for these manifolds. By a slam dunk on the knot with surgery coefficient $a/d$ into the 0 framed component, we obtain a knot $K_{0}$ with surgery coefficient d/a-d/a linked by the components $K_{1}$ and $K$ with surgery coefficients $b/c$ and abp/qab-p/q respectively. Together $K_{0}\cup K_{1}$ form a Hopf link where $K$ is a curve in the Heegaard torus between them that is isotopic to the meridian of $K_{0}$ and the longitude of $K_{1}$. Since ac+bd=1ac+bd=-1, a sequence of Rolfsen twists on $K_{0}$ and $K_{1}$ informed from a continued fraction expansion of d/a-d/a with partial fraction $b/c$ reduces both the surgery coefficients of $K_{0}$ and $K_{1}$ to $1/0$, twists $K$ to the $(a,b)$ curve on the Heegaard torus, and reduces its surgery coefficient to $p/q$. ∎
In [Is18] Issa classifies quasi-alternating Montesinos knots as follows.
Theorem 7.3 (Issa [Is18])
@@
    e=n1e=n-1 and $\frac{\beta_{i}}{\alpha_{i}}+\frac{\beta_{j}}{\alpha_{j}}<1$ for some $i\neq j$.
In addition, it is shown in [Is18, Corollary 1] that a Seifert fibered space is the double branched cover of a quasi-alternating link if and only if it is the double branched cover of a quasi-alternating Montesinos link. The Seifert fibered space $SFS[S2:(\alpha_{1},\beta_{1})\ldots(\alpha_{n},\beta_{n})]$ in Regina’s notation is the double branched cover of the Montesinos knot $M(0;\alpha_{i}/\beta_{i},\ldots,\alpha_{n}/\beta_{n})$ which is in general not in Issa’s standard form.
@@
Now by Theorem 7.2, $(p/q)$-surgery on $T_{a,b}$ is diffeomorphic to the (generalized) Seifert fibered space
@@
where $c$ and $d$ are any integers such that ac+bd=1ac+bd=-1. Since $a>b>0$ it follows that $d>0>c$. Furthermore the integers $c,d$ satisfying ac+bd=1ac+bd=-1 may be chosen so that $a>d>0$ and b>cb>-c, and hence $\frac{a}{d}>1$ and $\frac{b}{b+c}>1$.
Also, notice that if p/q=ab1p/q=ab-1 then the above manifold is a lens space, and thus the slope ab1ab-1 is quasi-alternating. From [Wa11] it follows that any slope $p/q\geq ab-1$ is quasi-alternating. Otherwise, we have p/q<ab1p/q<ab-1 so that $\frac{abq-p}{q}>1$ and Theorem 7.3 informs us that the Montesinos link $M\left(1;a/d,b/(b+c),ab-p/q\right)$ is quasi-alternating if and only if one of the following is fulfilled:
@@
(1) cannot occur because ad+bc=1ad+bc=-1. Either (2) or (3) occurs if and only if either $\frac{q}{abq-p}>\frac{-c}{b}$ or $\frac{q}{abq-p}>\frac{a-d}{a}$. Since abqp>q>0abq-p>q>0, $a>d>0$, and b>c>0b>-c>0, this is occurs if and only if $\frac{p}{q}>ab-\max\{\frac{a}{a-d},\frac{b}{-c}\}$. Thus, with [Is18, Corollary 1], this gives the claimed classification of quasi-alternating slopes once we set m=adm=a-d and n=cn=-c. ∎
Remark 7.4
For a knot $K$ in the 3-sphere, Owens and Strle [Owens_Strle12] define $m(K)$ to be the infimum of all $p/q>0$ such that $K(p/q)$ is the boundary of a smooth negative definite 4-manifold. It turns out that the lower bound in Theorem 7.1 coincides with $m(T_{a,b})$ with $a>b>1$. That is $p/q\in\mathbb{Q}$ is a quasi-alternating slope for $T_{a,b}$ if and only if $p/q>m(T_{a,b})$. Since every quasi-alternating surgery is the boundary of a smooth negative-definite 4-manifold, we have the inequality $m(T_{a,b})\leq ab-\max\left\{\frac{a}{m},\frac{b}{n}\right\}$. Equality readily follows by comparison with the formula for $m(T_{a,b})$ given in [Owens_Strle12, Theorem 2].
@@ 8. Formal L-space surgeries
In this section, we prove the results about formal L-space surgeries. As discussed in Remark 2.2 formal L-spaces can be seen as the $3$-manifold analogues of quasi-alternating links. First, we show that if the set of formal L-space surgery slopes is non-empty, then it is necessarily infinite.
@@
Let $K$ be a knot in $S^{3}$ and let $r=p/q\in\mathbb{Q}$ be a slope with $q\geq 1$. As illustrated in Figure 3, the manifold $K(p/q)$ is contained in a triad consisting of $Y_{0}=U(q/t)$, $Y_{1}=K(p/q)$, and $Y=K((p+q)/q)$, where p/q=nt/qp/q=n-t/q with $n\in\mathbb{Z}$ and an integer $t$ with $q>t\geq 0$. Recall that $K(p/q)$ has first homology isomorphic to $\mathbb{Z}_{p}$ and thus the triad $(Y_{0},Y_{1},Y)$ fulfills the homology condition in the definition of a formal L-space. Since every lens space is a formal L-space, we see that $K((p+q)/q)=K(p/q+1)$ is a formal L-space, whenever $K(p/q)$ is a formal L-space and $p/q>0$. Thus the theorem follows by induction. ∎
[…]
Figure 3. The triad obtained by performing 0, $1/0$ and 1-1 surgery on the knot $C$
When paired with the existence of quasi-alternating surgeries on asymmetric knots, this necessarily implies the existence of many formal L-spaces that are not obtained as the double branched cover of any link in $S^{3}$.
@@
Let $K$ be any asymmetric knot admitting an alternating or quasi-alternating surgery. For example, let $K$ be one of the asymmetric L-space knots in the SnapPy census or one of the Baker–Luecke asymmetric L-space knots. By Corollary 2.7, for all $r$ sufficiently large $K(r)$ is an asymmetric L-space. However, infinitely many of these surgeries are also formal L-spaces by Theorem 8.1. ∎
For integer surgeries yielding a formal L-space, Theorem 8.1 can be strengthened.
@@
where $a_{i}\geq 2$ for all $i$. We call $\ell$ the length of this continued fraction expansion. Note also that if $p/q>n$, then $a_{1}>n$. We prove the theorem by induction on $\ell$ and then by induction on the final coefficient $a_{\ell}$. The base case for this induction is the case $\ell=1$. In this case, $p/q$ is an integer and Theorem 8.1 gives the necessary result.
Thus suppose that $p/q>n$ has continued fraction expansion
@@
where $a_{i}\geq 2$ for all $i$ and $\ell>1$. Let $p_{0}/q_{0}$ be the rational number
$p_{0}/q_{0}=[a_{1},\dots,a_{\ell-2},a_{\ell-1}]^{-}.$
Since $a_{i}\geq 2$ for all $i$, we have that $p_{0}/q_{0}>a_{1}-1\geq n$ and the length of this continued fraction is $\ell-1$. Thus, we may inductively suppose that $K(p_{0}/q_{0})$ is a formal L-space. Let $p_{0}/q_{0}$ be the rational number
$p_{1}/q_{1}=[a_{1},\dots,a_{\ell-1},a_{\ell}-1]^{-}.$
Note that $p_{1}/q_{1}$ satisfies
$p_{1}/q_{1}=[a_{1},\dots,a_{k-1},a_{k}-1]^{-},$
where $k\leq\ell$ is maximal such that $a_{k}>2$. In particular, $p_{1}/q_{1}\geq a_{1}-1\geq n$ and by induction either on the length or final coefficient of the continued fraction expansion we may assume that $K(p_{1}/q_{1})$ is a formal L-space.
On the other hand, Figure 4 shows that the manifolds $K(p_{0}/q_{0})$, $K(p_{1}/q_{1})$ and $K(p/q)$ sit in a surgery triad. Furthermore standard continued fraction identities show that $\frac{p}{q}=\frac{p_{0}+p_{1}}{q_{1}+q_{0}}$. Thus we see that the homology groups of these manifolds satisfy the relation from the definition of a formal L-space. Since $K(p_{0}/q_{0})$ and $K(p_{1}/q_{1})$ are formal L-spaces, it follows that $K(p/q)$ is also a formal L-space. ∎
The fact that a Seifert fibered space is a formal L-space if and only if it is the double branched cover of a quasi alternating link [Is18] implies that for torus knots the set of formal L-space slopes coincides exactly with the set of quasi-alternating slopes.
@@
Surgery on a torus knot is a Seifert fibered space over $S^{2}$ (including the possibility of a lens space) or a connected sum of two lens spaces. From Theorem 7.2 it follows that $T_{a,b}(p/q)$ is a connected sum of two lens spaces if and only if $p/q=ab$ which is by Theorem 7.1 a quasi-alternating slope and thus also a formal L-space slope. In his classification of quasi-alternating Montesinos knots, Issa showed that a Seifert fibered space over $S^{2}$ is a formal L-space if and only if it is the double branched cover of a quasi-alternating link [Is18]. ∎
Finally we check that every non-trivial L-space knot has L-space surgery slopes that are not formal L-space surgeries.
@@
Like the double branched cover of a quasi-alternating link, a formal L-space is the boundary of a negative-definite manifold with $H_{1}=0$. If $Y_{0}$ and $Y_{1}$ are a pair of 3-manifolds appearing in a surgery triad for $Y$ as in the definition of a formal L-space, then it can be verified that one of the surgery cobordisms from $Y_{0}$ or $Y_{1}$ to $Y$ will be negative definite (see [Lidman2017, Section 2.2], for example). This allows one to build a negative-definite manifold inductively via 2-handle attachments to $B^{4}$. However, if $K(r)$ bounds a negative definite manifold with $H_{1}=0$, then $K(\lceil r\rceil)$ also bounds such a manifold, since there is a negative-definite 2-handle cobordism from $K(r)$ to $K(\lceil r\rceil)$ [Owens_Strle12, Lemma 2.6]. The work of Greene [Greene2015genusbounds] implies that we then have the bound
$2g(K)\leq\lceil r\rceil-\sqrt{\lceil r\rceil}$
which is equivalent to
$(\sqrt{2g(K)+1/4}+1/2)^{2}\leq\lceil r\rceil.$
Thus we obtain
$r>\lceil r\rceil-1\geq 2g(K)-\frac{1}{2}+\sqrt{2g(K)+1/4},$
which is equivalent to the claimed inequality. ∎
@@
Figure 4. The triad obtained by performing 0, $1/0$ and 1-1 surgery respectively on the knot $C$ yields $K(p_{0}/q_{0})$, $K(p_{1}/q_{1})$ and $K(p/q)$.
@@ 9. Observations and questions
The strategies from Section 3 together with the main result of [Twisting_QA] and [Watson_KH_sym2017] work well in practice to determine the quasi-alternating/thin status for large infinite subsets of slopes for a given knot. More concretely: For an asymmetric knot, every quasi-alternating/thin slope has to be one of the finitely many symmetry-exceptional-symmetric slopes, and thus one can evoke Strategies 3.3, 3.5, and 3.13. If an L-space knot $K$ is not asymmetric, then it is conjectured that it admits a single involution [L_space_periodic_conj]. In that case, we can use Strategy 3.9 to get all branching sets as tangle fillings $T(r)$ of the tangle exterior $T$. The main result of [Twisting_QA] gives a criterion to demonstrate that $T(r)$ is quasi-alternating for all $r$ in an interval of the form $a,\infty)$, while [Watson_KH_sym2017] can be used to show that infinitely $T(r)$ is Khovanov thick for all $r$ in an interval of the form $(-\infty,b)$. However, in general, $a$ and $b$ do not coincide.
@@
Consider the torus knot $T_{5,3}$. Theorem 7.1 shows that the quasi-alternating slopes are $(12+1/2,\infty)$. In Example 1.1 we have observed that the $(12+1/2)$-surgery on $T_{5,3}$ is the double branched cover of the knot $K11n50$ which is thin but not quasi-alternating [Greene_thin_QA]. If we apply [Watson_KH_sym2017] we get that all slopes smaller or equal to $12$ are thick. Thus the slopes in the interval $(12,12+1/2)$ remain of undetermined thick/thin status. However, experiments suggest that all surgeries in that interval are thin. For example: $K10n21(12+1/3)=DBC(K12n296)$ and $K10n21(12+1/4)=DBC(K13n2006)$ and both branching sets are thin but not quasi-alternating.
@@
What is the classification of the thin slopes of torus knots?
Most surgeries on torus knots are small Seifert fibered spaces [Moser1971elementary] and these arise as double branched covers of Montesinos links and Seifert links. Since the Seifert links with L-space double branched covers are also Montesinos links [ADE_link], this naturally leads to the question of thinness for Montesinos links.
@@
What is the classification of thin Montesinos links?
@@
Note that from the inductive definition of quasi-alternating links, it follows that there exists an algorithm that certifies in finite time if a knot is quasi-alternating. Thus Question 9.6 reduces to the existence question for an algorithm that certifies if link is not quasi-alternating.